OpenStudy (anonymous):

Help!! Win MEDAL!! Solve -2(3+6) = ? - 12

OpenStudy (uri):

PEDMAS.

OpenStudy (anonymous):

Need an answer! Win Medal!

OpenStudy (neer2890):

l.h.s -2(3+6)=-2*3+(-2)*6=-6-12 now relate this to right hand side and find the value of "?"

OpenStudy (anonymous):

-6?

OpenStudy (anonymous):

WIN MEDAL!

OpenStudy (anonymous):

Uri you too!

OpenStudy (anonymous):

Thanks for helping me you guys!

OpenStudy (neer2890):

you're welcome and just try to practice more problems to learn..:)